# Quantum-Adversary-Resilient Evidence Structures and Migration Strategies for Regulated AI Audit Trails

### v1

@@ 4.3 Security Theorems
    Otherwise, either $x^{\star}$ is a new input to the signing oracle or $\sigma^{\star}$ is a new signature on a previously queried $m^{\star}=H(x^{\star})$. In the former case, $(x^{\star},\sigma^{\star})$ constitutes a valid forgery against the EUF-CMA security of $\Sigma$; $\mathcal{B}$ outputs $(m^{\star},\sigma^{\star})$ as its forgery. In the latter case, if $\mathcal{B}$ did not query $H$ on $x^{\star}$, we can use standard techniques to extract a collision in $H$ from $\mathcal{A}$’s behavior in the QROM, contradicting collision resistance.
More formally, we construct $\mathcal{B}$ as a sequence of hybrids, gradually replacing oracle responses with simulated ones and bounding the statistical distance between experiments. The final hybrid yields either a signature forgery or a hash collision with probability non-negligibly related to $\epsilon(\lambda)$. Hence, if $\mathsf{Adv}^{\mathsf{qaudit}}_{\mathcal{E}}(\mathcal{A},\lambda)$ were non-negligible, one of $\mathsf{Adv}^{\mathrm{euf\text{-}cma}}_{\Sigma}(\lambda)$ or the collision-resistance advantage for $H$ would be non-negligible, contradicting our assumptions.
A fully rigorous proof in the QROM requires a careful treatment of quantum oracle simulation and programming; we omit these standard technical details here and refer to, e.g., [zhandry2012qrom, bonehzhandry2013qro, unruh2017pqsignatures] for the underlying techniques. ∎
A similar reduction shows Q-Binding: if an adversary can produce a single evidence item $\mathsf{Ev}^{\star}=(x^{\star},\sigma^{\star})$ that verifies as valid for two distinct events $e_{0},e_{1}$, then either $\mathsf{Fields}(e_{0})=\mathsf{Fields}(e_{1})$ (which we assume cannot happen for distinct events) or $\mathsf{Fields}(e_{0})\neq\mathsf{Fields}(e_{1})$, in which case the adversary has produced two different events with the same $x^{\star}$, contradicting the collision-resistance of the mapping $e\mapsto\mathsf{Fields}(e)$ (or of $H$ if $x$ is compressed further). This corresponds to the usual binding intuition: a single authenticator cannot be successfully opened as two different well-formed event descriptions except with negligible probability.
Non-equivocation can be shown under additional assumptions about the linkage mechanism (e.g., hash chains or Merkle trees) that prevent reordering and omission of evidence items, following techniques from transparency logs [ct]. We conjecture that Q-Non-Equivocation holds under standard assumptions when the linkage mechanism is instantiated with a collision-resistant hash function in the QROM; a full formalization and proof is left as future work.
@@ 5.1 Threat Scenarios: Harvest-Now, Forge-Later
In this scenario, the adversary does not need to compromise the system in real time: passive eavesdropping and later cryptanalysis suffice. This distinguishes audit migration from other settings (e.g., TLS) where the focus is on protecting present and future sessions rather than long-lived, highly replicable logs. In contrast to PQ-TLS migration, which focuses on protecting future sessions [alnahawi2024pqtls], audit trails must preserve the evidentiary value of past executions.
@@ 5.2 Migration Patterns
Auditors can then verify both the original classical evidence and its PQ wrapper: $\mathsf{Verify}^{\Sigma}(\mathsf{pk}^{\mathrm{pq}},H^{\prime}(\mathsf{Ev}_{i}),\tau_{i})$. The trust assumption is that the re-signing process is performed honestly and that the original log was unmodified at the time of re-signing. This process effectively extends the lifetime of legacy evidence by anchoring it to a PQ signature scheme in a single step.
@@ 5.3 Complexity and Storage Analysis
In practice, combinations of these patterns may be used: for example, Merkle-root anchoring for very old archival logs and full re-signing for recent, higher-value records. Recent system designs that combine PQC and blockchain for audit-grade file transfer [sola2025qrfile] follow a complementary pattern, anchoring events in an immutable ledger rather than constant-size evidence records.
@@ 6 Case Study: A Constant-Size Evidence System in Regulated AI
To illustrate how the above concepts translate into practice, we sketch a case study based on a constant-size evidence system implemented in an industrial platform for regulated AI workloads at Codebat Technologies Inc. The platform targets clinical and pharmaceutical workflows where each AI job (e.g., a risk prediction or trial cohort selection) is executed inside a hardened containerized environment and produces a compact evidence record.
@@ 6.1 Deployment Overview
In a typical deployment, the platform orchestrates AI jobs across a cluster of container-based trusted execution environments (TEEs). For each job, the runtime collects:
  • a hash of the container image (code and model);
  • digests of input datasets and configuration;
  • a platform measurement from the TEE (e.g., a measurement of the guest VM);
  • wall-clock timestamps and policy identifiers.
These fields are combined into a constant-size record $x=\mathsf{Fields}(e)$ and signed to produce an evidence item $\mathsf{Ev}=(x,\sigma)$, following the abstraction in Section 2.1. Evidence items are stored in an append-only log backed by a database; auditors can later retrieve evidence for any job and verify it against the public key of the platform.
In current deployments, the signature scheme is classical (e.g., ECDSA) while the rest of the design (constant-size layout, hash-based linkage) is agnostic to the choice of signature. The audit log may grow to $10^{6}$$10^{8}$ evidence records over several years, depending on workload volume.
@@ 6.2 Applying Migration Patterns
Once a PQ signature scheme is available in the platform’s cryptographic module, new jobs can produce hybrid evidence items $(x,\sigma^{\mathrm{class}},\sigma^{\mathrm{pq}})$ without changing the upstream workflow. The primary engineering task is to extend the evidence generator and verifier to handle an additional signature field. Because evidence records are already compact (e.g., $\approx 288$ bytes), adding a PQ signature of similar size roughly doubles the record size but keeps it under 1 KB, which is acceptable for many log storage systems.
@@
For recent evidence, especially those corresponding to higher-risk workflows (e.g., clinical trial analyses), the platform can re-sign individual records using a PQ signature key inside a TEE. For concreteness, we assume a PQ signature throughput of roughly 5,0005{,}000 signatures per second on commodity CPUs, in line with reported benchmarks for lattice-based signatures in TLS and PKI settings [raavi2025pqdsa]. Given $N_{\mathrm{recent}}$ such records, re-signing $10^{7}$ records would take on the order of 2,000 seconds ($\approx 34$ minutes) on a single core, or significantly less when parallelized. These back-of-the-envelope estimates suggest that re-signing at this scale is operationally feasible.
@@ 7 Related Work
Secure logging and transparency.
Classical secure logging and forward-secure audit log designs (e.g., [securelogging, forwardsecurelogging, haberstornetta1991, merkle1989certified]) focus on protecting logs against classical adversaries, often using hash chains and forward-secure keys. Transparency logs such as Certificate Transparency and Revocation Transparency [ct, transparency] add public append-only structures and gossip protocols to detect misbehavior. Our work builds on these foundations but explicitly targets quantum adversaries and constant-size evidence structures for regulated AI settings.
Post-quantum cryptography and logging.
Post-quantum cryptography has been extensively studied and standardized, e.g., [bernstein2009pqcbook, nistpqc, bonehshoup]. More recently, several works have explored quantum-resilient logging and auditing in domain-specific settings, such as distributed logs [somarapu2022qrlogs], cloud multimedia auditing [khan2025pqblockchainaudit], and telecom logs for 5G/6G [yusuf2025qrlogintegrity]. Survey data on PQC support in mainstream cryptographic libraries [ahmed2025pqclibs] indicates that hybrid and wrapper-based migration, as in our patterns, aligns well with how PQ algorithms are currently being integrated into software stacks. Our work is complementary: we focus on constant-size evidence structures and provide a unified treatment of quantum-adversary notions and migration patterns for regulated AI audit trails.
@@ 9 Conclusion
We have studied evidence structures for regulated AI audit trails in the presence of quantum adversaries. Building on an abstraction of constant-size evidence records, we introduced game-based notions of Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding, and analyzed a hash-and-sign instantiation in the quantum random-oracle model using post-quantum signature schemes. We then proposed migration patterns for deployed evidence logs—hybrid signatures, re-signing of legacy records, and Merkle-root anchoring—and examined their trade-offs.
@@
Future work includes extending our analysis to richer linkage mechanisms, integrating confidentiality and privacy requirements, and collaborating with regulators to align quantum-safe audit designs with evolving standards.

### v2

@@ 5.4 Security Theorems
    Case 2: $m^{\star}\in\mathcal{M}$. Then there exists a previously queried event $e$ with $H(\mathsf{Fields}(e))=m^{\star}=H(x^{\star})$. Since $(e^{\star},\mathsf{Ev}^{\star})\notin\mathcal{Q}$, either $\mathsf{Fields}(e^{\star})\neq\mathsf{Fields}(e)$ (yielding a collision in $H$, extracted by $\mathcal{B}_{2}$) or $\mathsf{Fields}(e^{\star})=\mathsf{Fields}(e)$ and $\sigma^{\star}\neq\sigma$ for the previously generated signature $\sigma$, in which case $(m^{\star},\sigma^{\star})$ is itself a valid forgery (a second valid signature on the same message produced without the signing key).
In either case, $\mathcal{A}$’s success translates to an advantage against $\Sigma$ or $H$:
$\epsilon(\lambda)\;\leq\;\mathsf{Adv}^{\mathrm{euf\text{-}cma}}_{\Sigma}(\mathcal{B}_{1},\lambda)\;+\;\mathsf{Adv}^{\mathrm{cr}}_{H}(\mathcal{B}_{2},\lambda),$
both of which are negligible by assumption. The simulation is perfect: all oracle responses seen by $\mathcal{A}$ are identically distributed to the real experiment, since $\mathcal{B}_{1}$ uses the genuine QROM oracle and the genuine signing oracle. Standard QROM simulation techniques (cf. [20, 5]) ensure that the reduction carries through even when $\mathcal{A}$ queries $H$ in quantum superposition. ∎
Theorem 2 (Q-Binding).
Under the same assumptions as Theorem 1, for any QPT adversary $\mathcal{A}$ against Q-Binding there exists a QPT adversary $\mathcal{B}$ against the collision resistance of $H$ such that
$\mathsf{Adv}^{\mathsf{qbind}}_{\mathcal{E}}(\mathcal{A},\lambda)\;\leq\;\mathsf{Adv}^{\mathrm{cr}}_{H}(\mathcal{B},\lambda).$
Proof.
Suppose $\mathcal{A}$ outputs $(e_{0},e_{1},\mathsf{Ev}^{\star})$ with $e_{0}\neq e_{1}$. Write $\mathsf{Ev}^{\star}=(x^{\star},\sigma^{\star})$ and assume $\mathsf{Verify}$ accepts $(e_{b},\mathsf{Ev}^{\star})$ for both $b\in\{0,1\}$.
By the verification algorithm (Section 2.1), $x^{\star}$ must equal both $\mathsf{Fields}(e_{0})$ and $\mathsf{Fields}(e_{1})$. Since $\mathsf{Fields}$ is deterministic, this requires $\mathsf{Fields}(e_{0})=\mathsf{Fields}(e_{1})$. Because each event description $e$ includes identifiers, timestamps, and artifact digests that are individually hashed to produce the fields of $\mathsf{Fields}(e)$ (see Table 1), having $\mathsf{Fields}(e_{0})=\mathsf{Fields}(e_{1})$ with $e_{0}\neq e_{1}$ implies that there exists at least one field index $j$ where the underlying data differs but the hash values coincide: $H(\phi_{j}(e_{0}))=H(\phi_{j}(e_{1}))$ with $\phi_{j}(e_{0})\neq\phi_{j}(e_{1})$, where $\phi_{j}$ extracts the $j$-th component from the event description. This constitutes a collision in $H$.
$\mathcal{B}$ simulates the full experiment for $\mathcal{A}$ and, upon receiving the output $(e_{0},e_{1},\mathsf{Ev}^{\star})$, iterates over the $k$ field positions to find the index $j$ where a collision occurs, outputting $(\phi_{j}(e_{0}),\phi_{j}(e_{1}))$. The extraction is deterministic, so the reduction is tight. ∎
Theorem 3 (Q-Non-Equivocation).
Let $\mathcal{E}$ be the hash-and-sign evidence structure with hash-chain linkage as defined in Section 5.3. For any QPT adversary $\mathcal{A}$ against Q-Non-Equivocation there exists a QPT adversary $\mathcal{B}$ against the collision resistance of $H$ such that
$\mathsf{Adv}^{\mathsf{qne}}_{\mathcal{E}}(\mathcal{A},\lambda)\;\leq\;\mathsf{Adv}^{\mathrm{cr}}_{H}(\mathcal{B},\lambda).$
Proof.
Suppose $\mathcal{A}$ outputs two logs $\mathcal{L}^{(0)}=\{(e^{(0)}_{i},\mathsf{Ev}^{(0)}_{i})\}_{i=1}^{n_{0}}$ and $\mathcal{L}^{(1)}=\{(e^{(1)}_{i},\mathsf{Ev}^{(1)}_{i})\}_{i=1}^{n_{1}}$ that both pass item-wise verification, produce matching chain tips $\ell^{(0)}_{n_{0}}=\ell^{(1)}_{n_{1}}$, and differ at some position.
We construct $\mathcal{B}$ as follows. $\mathcal{B}$ simulates the full experiment for $\mathcal{A}$, forwarding hash and signing queries as in Theorem 1. When $\mathcal{A}$ outputs the two logs, $\mathcal{B}$ computes both hash chains:
$\ell^{(b)}_{i}=H(\ell^{(b)}_{i-1}\;\|\;\mathsf{Ev}^{(b)}_{i}),\quad b\in\{0,1\},\;i=1,\dots,n_{b},$
starting from $\ell^{(b)}_{0}=0^{\lambda}$.
Case $n_{0}\neq n_{1}$. Without loss of generality, assume $n_{0}<n_{1}$. Since $\ell^{(0)}_{n_{0}}=\ell^{(1)}_{n_{1}}$ but $n_{0}<n_{1}$, we have $\ell^{(1)}_{n_{0}}\neq\ell^{(1)}_{n_{1}}$ in general (the chain continues for additional steps in log (1)). If $\ell^{(0)}_{n_{0}}=\ell^{(1)}_{n_{0}}$, then $\ell^{(1)}_{n_{0}}$ eventually maps to the same value $\ell^{(1)}_{n_{1}}=\ell^{(0)}_{n_{0}}=\ell^{(1)}_{n_{0}}$ through $n_{1}-n_{0}$ additional hash steps, producing a cycle in the hash chain that yields a collision in $H$. If $\ell^{(0)}_{n_{0}}\neq\ell^{(1)}_{n_{0}}$, then we are in the equal-length case for the prefix of length $n_{0}$ and proceed as below.
Case $n_{0}=n_{1}=n$. Both chains have the same length and produce the same tip: $\ell^{(0)}_{n}=\ell^{(1)}_{n}$. $\mathcal{B}$ walks backward from position $n$. Let $j^{\star}$ be the largest index $j\in\{1,\dots,n\}$ such that $\mathsf{Ev}^{(0)}_{j}\neq\mathsf{Ev}^{(1)}_{j}$ or $\ell^{(0)}_{j-1}\neq\ell^{(1)}_{j-1}$. Such an index exists because the logs differ at some position.
For all positions $i>j^{\star}$, we have $\mathsf{Ev}^{(0)}_{i}=\mathsf{Ev}^{(1)}_{i}$. By the maximality of $j^{\star}$, we also have $\ell^{(0)}_{j^{\star}}=\ell^{(1)}_{j^{\star}}$: this follows by downward induction from position $n$, since $\ell^{(b)}_{i}=H(\ell^{(b)}_{i-1}\|\mathsf{Ev}^{(b)}_{i})$ and for $i>j^{\star}$ both inputs agree, so the outputs agree. Therefore
$H\!\bigl(\ell^{(0)}_{j^{\star}-1}\;\|\;\mathsf{Ev}^{(0)}_{j^{\star}}\bigr)\;=\;\ell^{(0)}_{j^{\star}}\;=\;\ell^{(1)}_{j^{\star}}\;=\;H\!\bigl(\ell^{(1)}_{j^{\star}-1}\;\|\;\mathsf{Ev}^{(1)}_{j^{\star}}\bigr)$,
while by definition of $j^{\star}$ the inputs differ: $(\ell^{(0)}_{j^{\star}-1}\|\mathsf{Ev}^{(0)}_{j^{\star}})\neq(\ell^{(1)}_{j^{\star}-1}\|\mathsf{Ev}^{(1)}_{j^{\star}})$. $\mathcal{B}$ outputs this pair as a collision in $H$.
The collision extraction is deterministic (no guessing is required), so the reduction is tight: $\mathsf{Adv}^{\mathsf{qne}}_{\mathcal{E}}(\mathcal{A},\lambda)\leq\mathsf{Adv}^{\mathrm{cr}}_{H}(\mathcal{B},\lambda)$. ∎
@@ 6.1 Threat Scenarios: Harvest-Now, Forge-Later
In this scenario, the adversary does not need to compromise the system in real time: passive eavesdropping and later cryptanalysis suffice. This distinguishes audit migration from other settings (e.g., TLS) where the focus is on protecting present and future sessions rather than long-lived, highly replicable logs. In contrast to PQ-TLS migration, which focuses on protecting future sessions [2], audit trails must preserve the evidentiary value of past executions.
@@ 6.2 Migration Patterns
Auditors can then verify both the original classical evidence and its PQ wrapper by checking $\mathsf{Verify}^{\Sigma}(\mathsf{pk}^{\mathrm{pq}},H^{\prime}(\mathsf{Ev}_{i}),\tau_{i})$. The trust assumption is that the re-signing process is performed honestly and that the original log was unmodified at the time of re-signing. This process effectively extends the lifetime of legacy evidence by anchoring it to a PQ signature scheme in a single step.
@@ 6.3 Complexity and Storage Analysis
In practice, combinations of these patterns may be used: for example, Merkle-root anchoring for very old archival logs and full re-signing for recent, higher-value records. Recent system designs that combine PQC and blockchain for audit-grade file transfer [15] follow a complementary pattern, anchoring events in an immutable ledger rather than constant-size evidence records.
@@ 7 Case Study: A Constant-Size Evidence System in Regulated AI
To illustrate how the above concepts translate into practice, we present a case study based on a constant-size evidence system implemented in an industrial platform for regulated AI workloads at Codebat Technologies Inc. The platform targets clinical and pharmaceutical workflows where each AI job (e.g., a risk prediction or trial cohort selection) is executed inside a hardened containerized environment and produces a compact evidence record following the layout in Table 1.
@@ 7.1 Deployment Overview
In a typical deployment, the platform orchestrates AI jobs across a cluster of container-based trusted execution environments (TEEs). For each job, the runtime collects the attestation and provenance attributes described in Section 3 and produces an evidence record of 320–512 bytes (depending on signature scheme).
In current deployments, the signature scheme is classical (Ed25519) while the rest of the design (constant-size layout, hash-based linkage) is agnostic to the choice of signature. The audit log may grow to $10^{6}$$10^{8}$ evidence records over several years, depending on workload volume. With a baseline evidence generation throughput of $2.8\times 10^{5}$ records per second on a 16-core server (see Table 2), the system can handle sustained workloads of thousands of AI jobs per second while maintaining real-time evidence logging.
@@ 7.2 Applying Migration Patterns
Once a PQ signature scheme is available in the platform’s cryptographic module, new jobs can produce hybrid evidence items $(x,\sigma^{\mathrm{class}},\sigma^{\mathrm{pq}})$ without changing the upstream workflow. The primary engineering task is to extend the evidence generator and verifier to handle an additional signature field. Because evidence records are already compact (320–512 bytes as in Table 1), adding a PQ signature (e.g., Dilithium at $\approx 2{,}400$ bytes or SPHINCS+ at $\approx 7{,}800$ bytes) increases the record size to 3–8 KB, which remains acceptable for most log storage systems given the strong long-term security guarantees.
Post-quantum signature performance.
The choice of PQ signature scheme directly affects evidence record size and migration throughput. Table 4 compares the classical baseline (Ed25519) with three PQ candidates standardized by NIST [12]: ML-DSA (FIPS 204, lattice-based, formerly CRYSTALS-Dilithium) at two security levels, and SLH-DSA (FIPS 205, hash-based, formerly SPHINCS+) in its small-signature variant. Performance figures are single-core reference-implementation estimates on commodity server hardware (3.0 GHz x86-64), drawn from recent benchmarks [13, 17].
Table 4: Post-quantum signature scheme comparison for evidence record migration. Evidence size = $k\times 32$ bytes (fields) + signature size. Performance figures are approximate single-core values from reference implementations [13].
Scheme $|\sigma|$ (B) $|\mathsf{pk}|$ (B) Sign/s Verify/s Ev size
Ed25519 (baseline) 64 32 35,000 12,000 320 B
ML-DSA-65 (Level 3) 3,309 1,952 5,000 12,000 3.6 KB
ML-DSA-87 (Level 5) 4,627 2,592 3,000 8,000 4.9 KB
SLH-DSA-128s (Level 1) 7,856 32 5 200 8.1 KB
ML-DSA-65 offers the best balance for evidence migration: its signing throughput of $\approx 5{,}000$ operations per second per core is sufficient for re-signing large corpora, while the $3.3$ KB signature keeps evidence records under $4$ KB. SLH-DSA provides the most conservative security assumptions (pure hash-based) but is significantly slower for signing, making it better suited for anchoring small batches rather than re-signing individual records.
@@
For recent evidence, especially those corresponding to higher-risk workflows (e.g., clinical trial analyses), the platform can re-sign individual records using a PQ signature key inside a TEE. Using ML-DSA-65 at $\approx 5{,}000$ signatures per second per core, re-signing $10^{7}$ records would take on the order of 2,000 seconds ($\approx 34$ minutes) on a single core, or under 3 minutes when parallelized across 16 cores. These estimates suggest that re-signing at this scale is operationally feasible, even accounting for the $\approx 7\times$ slowdown compared to classical Ed25519 signatures.
@@ 8 Related Work
Forensics and audit logging.
A rich body of work studies secure logging, accountability mechanisms, and digital forensics for regulated environments. Classical secure audit logs [14] and forward-secure logging schemes [3] focus on protecting log integrity against adversaries who may compromise the logging system after the fact. Foundational work on cryptographic timestamping [6] and authenticated data structures [11] provides building blocks for tamper-evident logs. More recent work explores accountability and audit trails in cloud and distributed settings, including blockchain-based approaches for immutable record-keeping [15, 8]. Our work complements these approaches by treating constant-size evidence structures as a first-class cryptographic object, with explicit security definitions against quantum adversaries and migration patterns tailored to already-deployed regulated AI systems.
Table 5 provides a systematic comparison of our work with representative systems from each category.
Table 5: Feature comparison with related work on secure logging and post-quantum audit systems. Symbols: ✓ = supported, $\circ$ = partial, – = not addressed.
Constant-size
PQ signatures
QROM analysis
Formal proofs
Migration
CT [10] $\circ$
Somarapu [16]
Khan et al. [8]
Sola-Thomas [15]
Yusuf [19]
This work
Transparency logs.
Transparency logs such as Certificate Transparency [10] and Revocation Transparency [9] add public append-only structures and gossip protocols to detect misbehavior by certificate authorities and other trusted parties. These designs have inspired similar transparency mechanisms in other domains, including software supply chain and binary transparency. Our evidence structures share the goal of providing tamper-evident audit trails, but focus on per-execution attestation in regulated AI settings rather than public certificate ecosystems.
Post-quantum cryptography.
Post-quantum cryptography has been extensively studied and is now being standardized through efforts such as the NIST PQC project [4, 12]. Signature schemes based on lattices, hash functions, and other post-quantum assumptions are reaching deployment maturity [17, 13], with growing support in mainstream cryptographic libraries [1]. Several works have begun to explore quantum-resilient logging and auditing in domain-specific settings, such as distributed log storage [16], cloud multimedia auditing [8], telecom logs for 5G/6G [19], and post-quantum TLS migration [2]. Our work provides a unified treatment of quantum-adversary security notions for constant-size evidence structures and concrete migration patterns for regulated AI audit trails.
@@ 9 Discussion and Open Problems
Richer linkage mechanisms.
Our Q-Non-Equivocation proof (Theorem 3) covers hash-chain linkage. Extending this to richer structures such as Merkle-based transparency logs, skip-list constructions, or DAG-based provenance graphs under quantum adversaries is an interesting direction. Such extensions would enable quantum-safe non-equivocation guarantees for a broader class of audit trail architectures beyond linear chains.
@@
Data and Code Availability
The prototype implementation used for the microbenchmarks in Table 2 and for the case study in Section 7 is part of an industrial platform developed by Codebat Technologies Inc. and is currently involved in ongoing patent applications. Due to commercial and contractual constraints, the source code and deployment configuration cannot be released at this time. The algorithms, system architecture, and parameter choices described in Sections 3 and 6 are sufficient to reproduce the reported complexity trends and to re-implement compatible evidence generators and migration services on comparable hardware. If and when the relevant patents are granted and contractual constraints permit, we plan to release a simplified reference implementation for research and interoperability purposes.
@@ 10 Conclusion
We have studied evidence structures for regulated AI audit trails in the presence of quantum adversaries. Building on an abstraction of constant-size evidence records, we introduced game-based notions of Q-Audit Integrity, Q-Non-Equivocation, and Q-Binding, and proved via tight reductions that a hash-and-sign instantiation in the quantum random-oracle model satisfies all three properties under standard assumptions (collision-resistant hashing and quantum EUF-CMA signatures). We then proposed migration patterns for deployed evidence logs—hybrid signatures, re-signing of legacy records, and Merkle-root anchoring—and examined their trade-offs with concrete post-quantum signature benchmarks.
@@
Future work includes extending our analysis to richer linkage mechanisms (e.g., Merkle-based transparency logs and DAG-based provenance graphs), integrating confidentiality and privacy requirements via post-quantum encryption, developing composable security frameworks for evidence structures embedded in larger protocols, and collaborating with regulators to align quantum-safe audit designs with evolving standards.